OpenStudy (anonymous):

check my answers To calculate the perimeter of a square, which formula(s) can be used?

OpenStudy (anonymous):

P = s2 P = 4 • s <-- my answer P = s + s + s P = s + s + s + s <- my answer

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes the two you marked are equivalent. P = 4s is the simplified form